### v2

# ChIRAAG: ChatGPT Informed Rapid and Automated Assertion Generation

Abstract
System Verilog Assertion (SVA) formulation- a critical yet complex task is a prerequisite in the Formal Property Verification (FPV) process. Traditionally, SVA formulation involves expert-driven interpretation of specifications, which is time-consuming and prone to human error. However, LLM-informed automatic assertion generation is gaining interest. We designed a novel framework called ChIRAAG, based on OpenAI GPT4, to generate SVA assertions from natural language specifications. ChIRAAG constitutes the systematic breakdown of design specifications into a standardized format, further generating assertions from formatted specifications using LLM. Furthermore, we developed testbenches to verify/validate the LLM-generated assertions. Automatic feedback of log files from the simulation tool to the LLM ensures that the framework can generate correct SVAs automatically. Only 33% of LLM-generated raw assertions had errors. Our results on OpenTitan designs shows that LLMs can streamline and assist engineers in the assertion generation process, reshaping verification workflows.
Index Terms:
Assertion Generation, LLM, Formal Verification
I Introduction
Large Language Models (LLMs) have evolved to assist humans in solving complex problems. LLMs create numerous possibilities for transformative impact across different domains, including education [1], medicine [2], and finance [3]. The hardware design industry has used LLMs to execute steps of Electronic Design Automation (EDA) at a faster rate[4].
Pre-silicon bug detection is a crucial step in designing functionally correct hardware. Security vulnerabilities and bugs are identified during the pre-silicon phase via Formal Property Verification (FPV). As per the functional verification study conducted by Wilson Research Group in 2020 [5], over 50% of the overall development resources and expenditures in Application-Specific Integrated Circuit (ASIC) and Field-Programmable Gate Array (FPGA) systems were allocated to FPV. FPV is a systematic and rigorous process of using formal methods to verify whether a system adheres to its specifications. FPV ensures that the system meets specified properties under all possible scenarios [6].
Assertions Based Verification (ABV) are used to verify the properties in over 50% of the designs  [5]. ABV validates the correct design implementation by integrating safety and liveness properties, expressed as assertions, throughout the development process. Safety property states that any path (sequence of events or states) violating the property has a finite prefix (initial segment of the path) such that every extension of the prefix violates the property. For example, “whenever a signal request is asserted, signal acknowledgment is asserted within three cycles” is a safety property [6]. While a liveness property states that any finite path (sequence of events or states) can be extended to a path satisfying the property [6]. For example, “Every request is eventually acknowledged” is a liveness property. However, as the scale of design and verification requirements increases, so does the number of assertions, which can reach hundreds, making it too difficult to be proven formally or falsified within the constraints of available resources manually [7].
LLMs could perform ABV quickly by automatically generating the correct assertions from the natural language specification that holds in simulation. The existing methods have relied on generating SVA from Verilog code [8], comments, and assertion-based examples [9]. None of them could generate SVA assertions directly from specification. Also, they still require potential intervention and expertise from the experts. This is the motivation for this study to formulate a novel framework that reduces human effort and time in ABV. This work aims to analyze how LLMs can assist ABV by generating assertions automatically from the natural language specifications. In our framework, manual intervention is needed only when there is a potential bug in the design implementation. We seek to answer two key questions through this work: RQ1: How good are LLMs in generating System Verilog Assertion (SVA) from natural language specification? RQ2: How functionally correct are the generated assertions within the design and verification framework?.
Our contributions are fourfold.
  • A framework to study the use of LLM in assertion generation for ABV.
  • A prompting method to provide relevant information to an LLM to generate SVA.
  • An analysis of the performance of LLMs in generating SVA for OpenTitan designs [10].
  • Recommendations of LLMs for assertion generation.
II Background and related works
II-A Overview of LLM
LLMs are Transformer Language models where the construction of the weight parameters involves training models using an extensive volume of text data acquired from diverse sources such as various websites, web-books, and PDFs. Based on their functionality, LLMs can be divided into two: General Domain LLMs (gd-LLMs) and Domain-specific LLMs (ds-LLMs) [11]. gd-LLMs are trained on diverse resources from various domains and sources, making them applicable to a wide range of queries. Examples of gd-LLMs include Llama 2 [12], and GPT4 [13]. While ds-LLMs focus on a particular realm of interest. Examples include MedicalGPT [14] and FoodGPT [15]. LLMs are expensive to train on large datasets, so fine-tuning pre-trained LLMs is becoming important to simplify training. Fine-tuning of an LLM is a supervised process to adjust the model parameters based on the task [16]. Code Llama [17] by Meta and StarCoder [18] by Hugging Face are some examples.
II-B LLM-Driven Automatic Assertion Generation
LLMs have leading-edge capabilities in diverse tasks, including source code generation [19], automated testing [20], and program repair [21, 22]. Automatic assertion generation is gaining popularity in recent times [8]. This study showcased the capability of LLM to generate assertions solely from RTL code without additional specifications. Despite initial syntactic and semantic errors, the authors observed that, with proper guidance, LLMs consistently produce correct and complete assertions. The work in [9] uses comments about the security assertions to generate assertions. They assessed the effectiveness of LLMs in accelerating property verification.They generated 75,600 assertions across diverse conditions and varying levels of context and found that LLMs can generate correct assertions at an average rate of 4.53%. They also mentioned that detailed comments and context proved to be particularly effective in generating correct assertions. The work in [23] uses LLMs in security assertion generation. They perform formal property verification to identify vulnerabilities before developing the Design Under Test (DUT). The generated assertions at the pre-silicon level can help monitor and mend security vulnerabilities during the initial phases of design.
In contrast, we have exclusively used the design specification in natural language as an input to LLM to generate SV assertion. The design specification operates at an abstract level, serving as a representation of the design, in contrast to directly providing RTL code for assertion generation. Since the objective of assertions is to identify the bug in the implementation. Therefore, it is not desirable to obtain the assertions from implementation. Moreover, creating assertions from a design specification poses a significant challenge in the absence of a dedicated and specialized framework. Therefore, we proposed a novel framework that formats the input specification to the LLM and further reviews, analyzes, and mends the LLM-generated assertions, if required. We have mainly focused on generating functional assertions rather than generating security assertions. None of the above-discussed works generate assertions merely using specification.
II-C The Three C’s in ABV
ABV uses assertions, which are formal statements or conditions that must always hold in a design for the proper functioning of logic. A primary objective of an assertion is to ensure consistency in the context of the designer’s intent and the actual realization of the design [24]. “Consistency” denotes no conflict or contradiction between the formal properties/specifications that could lead to ambiguous or erroneous results. “Correctness” refers to the properties where it assures that the assertions correctly represent the intended behavior of the design. “Completeness” states that the assertions should cover all related conditions and edge cases within the design. The verification process shouldn’t miss critical aspects of system behavior. For example, in a simple digital counter,
assert @(posedge clk_i) (count == previous_count + 1);
represent a correct assertion that ensures the counter increments correctly with each clock cycle by checking that the current count equals the previous count plus one. While,
assert @(posedge clk_i) (enable_signal == (start_signal && !stop_signal));
represents a consistent assertion that ensures control signals are handled consistently. This case checks that the enable signal is true only when the start signal is true and the stop signal is false, preventing contradictory combinations of control signals in the digital counter circuit. And,
assert @(posedge clk_i) (count >= 0 && count <= MAX_COUNT);
represents a completeness assertion that ensures that the counter’s value remains within the valid range. It checks that the count is always $\geq 0$ and does not exceed the maximum count value (MAX_COUNT). The main motivation of our work is to assist the engineer in generating the assertions using LLMs directly from the specifications. In this work, we mainly focused on the correctness of the assertions that are generated by the LLMs.
III ChIRAAG: Our SVA Generation Framework
The automatic assertion generation with ChIRAAG involves the compilation of two significant stages. Algorithm 1 shows the steps of our proposed ChIRAAG framework. While the stages are briefly described below:
III-A Formatting of specification
The formatting of design specifications includes representing the design context with specific labels informing about its general definitions, signals, functional parameters, etc. The formatting involves extracting the important information from the design specification, $\mathcal{D}$, and later converting the extracted information, $UnFmt$, into a Javascript Object Notation (JSON) format, $\mathcal{J}$. This operation can be symbolically represented as $Extraction:\mathcal{D}\rightarrow UnFmt$ followed by $JSON:UnFmt\rightarrow\mathcal{J}$, as represented in Line 7-10 of Algorithm 1. The $JSON$ function comprises several sub-functions, each responsible for structuring a specific aspect of the extracted information into JSON format, $\mathcal{J}$, where $\mathcal{J}$ includes {{\$\mathcal{J}_{int}$, $\mathcal{J}_{s\_ov}$, $\mathcal{J}_{df}$, $\mathcal{J}_{par}$, $\mathcal{J}_{fn\_r}$, $\mathcal{J}_{tm\_r}$, $\mathcal{J}_{ext}$}}\}}. The components in Table I show the formatting functions with their respective labels.
TABLE I: Formatting of specification
Label Description
Introduction ${\mathcal{J}_{int}}\leftarrow JSON_{int}$($UnFmt$)
System_overview ${\mathcal{J}_{s\_ov}}\leftarrow JSON_{s\_ov}$ ( $UnFmt$)
Definitions $\mathcal{J}_{df}\leftarrow JSON_{df}$ ($UnFmt$)
Parameters $\mathcal{J}_{par}\leftarrow JSON_{par}$ ($UnFmt$)
Functional_requirements $\mathcal{J}_{fn\_r}\leftarrow JSON_{fn\_r}$ ($UnFmt$)
Timing_requirements $\mathcal{J}_{tm\_r}\leftarrow JSON_{tm\_r}$ ($UnFmt$)
Extra_info $\mathcal{J}_{ext}\leftarrow JSON_{ext}$ ($UnFmt$)
III-B Automatic Assertion Generation
This is the second stage of ChIRAAG, which mainly involves the generation of the SV assertions using the JSON formatted specifications. The formatted specifications of the designs are provided as input to the LLM through our framework by making a call to OpenAI Application Programming Interface (API). This generates raw SV assertions as stated in Line 13 of Algorithm 1. The generated assertions are further verified for correctness through simulation runs utilizing testbenches in a simulation environment, producing simulation logs. This step gives rise to four possible cases that can eventually occur. They are:
No Error Message
When the generated assertion, $A_{n}$, passes all the test cases, i.e., the log file, $\mathcal{L}$, contains no errors, then $A_{n}$ is taken as correct output assertions suite, $S$. Lines 15-16 of Algorithm 1 shows that when $\mathcal{L}$ contains no error message, the while loop gets terminated, and then the ‘OUTPUT GENERATION’ procedure, in Lines 34-37, gets executed.
Syntax Error
Some LLM-generated assertions have Syntax Errors (SE). LLM is automatically instructed about the syntax error and also to refine the assertion. It is able to correct the syntax error within a few iterations. Lines 17-18 of Algorithm 1 check if the log file, $\mathcal{L}$, contains any SE, then the function ‘SyntaxRectifier’ gets executed. ‘SyntaxRectifier’ takes the assertions, $A_{n}$, and log file, $\mathcal{L}$, as input, and a syntax-based prompt as a heading is provided automatically to the LLM using the file-handling process. This process generates the refined assertions, $\overline{A_{n}}$.
Simulation Failure
The assertions generated by LLM might fail in simulation run because of specific errors. Two such simulation errors we have encountered are timing errors and missing signals. Lines 20-21 of Algorithm 1 check if the log file, $\mathcal{L}$, has any timing or missing signal error, then similar to SE, here the function ‘SimulationRectifier’ takes the assertions, $A_{n}$, and the log file, $\mathcal{L}$, as input and further a simulation-based prompt as a heading is provided automatically to LLM using the file-handling process. This leads to the generation of the rectified assertions, $\overline{A_{n}}$.
Implementation Bug
Sometimes, it might be possible that even the generated assertions are correct, but the implementation might be buggy. In such a scenario, the testbench may give out a wrong output, resulting in testcase failure. In such cases, manual inspection of design implementation is to be done as shown in Lines 23-24 of Algorithm 1.
The assertions may not pass the test cases in a single run. Therefore, we iteratively run the error-checking process multiple times. Line 14 of Algorithm 1 shows that the while loop runs until $n$ is less than the upper bound $\mathcal{T}$. If $n$ gets equal to $\mathcal{T}$, and there are still error messages in the log file, $\mathcal{L}$, we manually examine the design implementation for potential bugs, as stated in Lines 29-33. Also, the newly generated assertions, $\overline{A_{n}}$, is asserted to $A_{n}$, as stated in Lines 19 and 22 of the Algorithm 1, to follow the further error-checking execution steps in the loop. If there will be no error message in the log file, $\mathcal{L}$, and $n$ is less than $\mathcal{T}$, we execute the Lines 34-37 of Algorithm 1 as discussed in `$No\text{ }Error\text{ }Message^{\prime}$ case to get our final SV assertions, $S$. The proposed ChIRAAG framework is illustrated in the Fig. 1.
Figure 1: ChIRAAG-SVA Generation Framework
Algorithm 1 ChIRAAG: Automated SVA Generation Framework
1:Input:
2:      $\mathcal{D}$: Design specification document
3:      $\mathcal{L}$: Log file of simulation tool.
4:      $\mathcal{T}$: Upper bound on the number of LLM interaction prompts
5:Output:
6:      $S$: Suite of System Verilog Assertions
7:procedure Specification Extraction and Formatting
8:    $UnFmt\leftarrow Extraction(\mathcal{D})$
9:    $\mathcal{J}\leftarrow\text{JSON}(UnFmt)$
10:end procedure
11:procedure LLM Interaction for SVA Generation
12:    Initialize $n\leftarrow 0\text{ and }\mathcal{L}\leftarrow\{\}$
13:    $A_{n}\leftarrow\text{LLM}(\mathcal{J},\mathcal{L})$
14:    while $n<\mathcal{T}$ do
15:         if no errors in  of $\mathcal{L}\text{ of }A_{n}$ then
16:             break
17:         else if syntactic errors in  of $\mathcal{L}\text{ of }A_{n}$ then
18:             $\overline{A}_{n}\leftarrow\text{SyntaxRectifier}(A_{n},\mathcal{L})$
19:             $A_{n}\leftarrow\overline{A_{n}}$
20:         else if timing or missing signal error in  of $\mathcal{L}\text{ of }A_{n}$ then
21:             $\overline{A}_{n}\leftarrow\text{SimulationRectifier}(A_{n},\mathcal{L})$
22:             $A_{n}\leftarrow\overline{A_{n}}$
23:         else if wrong output due to testcase failure then
24:             Manual inspection of design implementation
25:             break
26:         end if
27:         $n\leftarrow n+1$
28:    end while
29:end procedure
30:procedure Design Implementation Examination
31:    if $n==\mathcal{T}$ and error(s) in $\mathcal{L}$ then
32:         Examine design implementation for bugs
33:    end if
34:end procedure
35:procedure Output Generation
36:    $S\leftarrow{A}_{n}$
37:    return $S$
38:end procedure
IV Case Study: RV Timer
Our proposed Algorithm 1 on SVA generation is assessed using multiple designs. However, we are taking a particular design, RV Timer, to mention our findings.
The RV Timer has clk_fixed as the fixed clock input and prescaler that divides the input clock frequency to a lower frequency. This lower frequency is later used by the timer. The timer’s count is incremented based on timer ticks generated, which is determined by the prescaler value. The generated ticks are sent to multiple hardware threads. A step value serves as an input to the timer, determining the increment step for each tick. The mtime register is the main timer counter that gets incremented with each tick and holds the current tick value. The comparison to check if a timer has reached a specific time value is performed using the mtimecmp register. The output signal expired is triggered when the mtime value is greater than or equal to the mtimecmp register value. The block diagram of the RV Timer is shown in Fig . 2
Figure 2: Block Diagram of RV TIMER
The unstructured design specification of RV Timer is initially transformed to structured JSON format and presented as input to the LLM to generate the SV assertions. Listing 1 provides a part of the structured format specification of RV Timer.
Listing 1: RV Timer Module Specification
"Introduction": {
"Purpose": "This document specifies the operational, functional, and timing requirements for the RV Timer module, which is designed to provide timing functions for the system.",
"System_Overview": {
"Architecture_Summary": "The RV Timer module includes tick generators, counters, comparators, and is capable of raising interrupts based on configurable timing events.",
"Definitions": {
"Signals": [
"Name": "clk_fixed",
"Type": "Input",
"Width": "As per system configuration",
"Description": "Fixed clock input for the timer."
"Parameters": [
"Name": "prescaler",
"Type": "Configuration",
"Description": "Prescaler value for the fixed clock to generate the timer tick."
"Functional_Requirements": {
The LLM initially generated 8 functional assertions. We further evaluated the correctness of the raw assertions using our testbench and simulator. Of the 8 assertions generated, 5 were correct in their raw format. This shows the potential of our proposed framework ChIRAAG. The remaining 3 raw assertions were rectified automatically using the steps of Algorithm 1. This process generated an extra assertion, which is essential to cover all the design aspects, increasing the count of generated assertions to 9. Some of the initial assertions of RV Timer generated by the LLM are:
Assertion 1:
// Assertion to check if tick_count resets to 0 on reset
property p_reset_tick_count;
@(posedge clk_i) (!rst_ni) |-> (tick_count == 12’h0);
end property
assert property (p_reset_tick_count);
Assertion 2:
// Assertion to check if tick_count increments correctly
property p_tick_count_increment;
@(posedge clk_i) (active && (tick_count < prescaler)) |-> (tick_count == \$past(tick_count) + 1);
end property
assert property (p_tick_count_increment);
The Assertion 1 is a correct raw assertion that passes all the tests. It ensures that ‘tick_count’ resets to 0 when reset is asserted, i.e., when ‘rst_ni’ or the active low reset signal is not active. While the majority of assertions are correct in their raw format, a few failed to pass certain test cases due to the presence of syntactic error or simulation error. As in Assertion 2, we observed that a missing signal prevents passing all the test cases as shown in Fig. 3. The LLM understood the missing signal rst_ni from the learned information and included it in the generated assertion. The ‘SimulationRectifier’ function shown in Algorithm 1 is called upon to rectify the Assertion 2. The LLM generated the refined assertion and modified it into two distinct assertions Modified Assertion 2a and Modified Assertion 2b.
Figure 3: Failure log of Missing signal
Modified Assertion 2a:
// Assertion for tick_count increment
property tick_count_increment;
@(posedge clk_i)
(rst_ni && active && (tick_count < prescaler)) |=> (tick_count == \$past(tick_count) + 1);
endproperty
assert property (tick_count_increment);
Modified Assertion 2b:
// Assertion for tick_count reset on reset deassertion
property tick_count_reset_on_reset_deassert;
@(posedge clk_i)
(!\$past(rst_ni, 1) && rst_ni) -> (tick_count == 0);
endproperty
assert property (tick_count_reset_on_reset_deassert);
The Modified Assertion 2a ensures that as long as the system is not in reset, is active, and the tick count is below a set threshold, the tick count will increment by 1 in each clock cycle. The Modified Assertion 2b verifies that when the system transitions from a reset state to a non-reset state, the tick count is reset to 0, ensuring proper initialization or re-initialization of the system’s tick counter. Together, Modified Assertion 2a and Modified Assertion 2b form a comprehensive set of checks for the tick_count signal, covering both its incremental behavior during normal operation and its initialization behavior upon reset deassertion. These assertions ensure the tick_count signal behaves correctly across different operational states of the system, contributing to the overall reliability of the hardware functioning.
V Experimental Results and Discussion
Our novel framework ChIRAAG, was evaluated using the designs provided in the open-source project OpenTitan [10]. We have performed our experiments on seven designs of the OpenTitan project, which includes RV Timer, PattGen, GPIO, HMAC, ROM Ctrl, sram_ctrl and adc_ctrl.
We have utilized OpenAI’s GPT-4 turbo [13] LLM as an automatic assertion generation tool for our ChIRAAG framework. This model has a capacity of 1280000 tokens of context window and training data up to December 2023. We collected the design specification and related files for each design and manually structured the data in a specific format with distinctive labels, and corresponding details for each design.
ChIRAAG generated the correct assertion using LLM within a few automatic prompting processes for each design. Table II shows the performance report for LLM-based assertion generation. The first column, “Module” shows different designs we have taken. The third column shows the number of SV assertions generated by LLM for each design, and the second column, “OT Assert.”, shows the number of assertions provided by the RTL in OpenTitan. “Iterations to Conversion” denotes the number of iterations required to re-prompt the LLM to generate the correct assertions that satisfy all the test cases. “VCS Sim. Time” shows the time taken by the simulation process. While the “ChIRAAG Time” is the SVA generation time taken by our framework.
ChIRAAG successfully generated all of the OpenTitan assertions for each design. Interestingly, ChIRAAG generates more assertions than that was provided in OpenTitan. For example, it generates 12 assertions for HMAC while only four assertions are provided in OpenTitan. Importantly, ChIRAAG-generated assertions are important and cover some aspects of the design that were missed by the OpenTitan assertions. For example, Assertion 3:
Assertion 3:
assert @(posedge clk) disable iff (rst_n)
(msg_fifo_req && hmac_en) |-> ##1 reg_hash_start)
else \$fatal("Assertion failed");
is crucial to check the HMAC process start property. The HMAC process should start whenever there is a request in the message FIFO and the HMAC signal is enabled. That is, if ‘msg_fifo_req’ and ‘hmac_en’ are true, then the signal ‘req_hash_start’ must be true on the next rising edge of the clock unless a reset (‘rst_n’) is active. If this condition is violated, a fatal error message is triggered. While in the ROM Controller, the additional important SVA that ChIRAAG generated besides the ones that are provided in OpenTitan assertions are:
Assertion 4:
assert @(posedge clk) (checker_done && (q !== Checking && q !== Done’)) |-> ##1 (q $=$ Invalid’);
Assertion 5:
assert @(posedge clk) (counter_done |-> (q !== Reading Low && q_next !== Invalid’)
Assertion 6:
assert @(posedge clk) (((q !== Invalid’) && rst_ni && kmac_rom_last_o) |-> ##1 kmac_rom_vld_o)
These assertions verify some of the essential properties: Assertion 4 checks if ‘checker_done’ is true when the system state (‘q’) is not in the ‘Checking’ or ‘Done’ states. If the condition is violated, the assertion requires that in the next clock cycle, the state ‘q’ must be set to invalid. Assertion 5 states that if the counter_done signal is true, then the state ‘q’ is either not ‘Reading Low’ in the current state or does not become ‘Invalid’ in the next state. While, Assertion 6 states that if the current state ‘q’ is not ‘Invalid’ and both the reset condition ‘rst_ni’ and the signal ‘kmac_rom_last_o’ are true, then in the next clock cycle, the signal ‘kmac_rom_vld_o’ should be true.
TABLE II: Performance Report
Module OT Assert. LLM Assert. Iteration to Conversion VCS Sim. Time ChIRAAG Time
RV Timer 0 9 9 80ns 7.34s
PattGen 0 7 5 120ns 8.35s
GPIO 0 6 10 105ns 9.24s
HMAC 4 12 15 180ns 8.29s
ROM Ctrl 6 14 16 110ns 9.42s
sram_ctrl 0 11 3 160ns 8.34s
adc_ctrl 5 7 5 95ns 7.29s
Also, generating assertions traditionally (manually) is a hectic and time-consuming process. Manually for each design, it would take many hours. Our proposed framework ChIRAAG, does the same work in less than 15 seconds for each design. This shows the effectiveness of ChIRAAG in a faster assertion generation process than the traditional method.
While ChIRAAG can rectify syntactic and simulation errors, it can also detect potential bugs in the implementation. To demonstrate such a scenario, we have considered a buggy implementation of a 4-bit full-adder, where xor is replaced by an or operation. Initially, formatted specifications were fed to ChIRAAG to generate the SVA. ChIRAAG correctly generated two assertions listed as Assertion 7 and Assertion 8, for checking the carry-out generated and the other for checking the sum bits generated.
Assertion 7:
//Assertion to check if the carry-out (C4) is generated correctly for the addition
property prop_carry_out;
@(posedge C0)
C4 |-> ((A[3] & B[3]) | (C3 & (A[3] ^ B[3])));
endproperty
assert property (prop_carry_out);
Assertion 8:
//Assertion to check if the sum bits (S) are correct
property prop_sum_bits;
@(posedge C0)
S[0] |-> (A[0] ^ B[0] ^ C0) &&
S[1] |-> (A[1] ^ B[1] ^ C1) &&
S[2] |-> (A[2] ^ B[2] ^ C2) &&
S[3] |-> (A[3] ^ B[3] ^ C3);
endproperty
assert property(prop_sum_bits);
Even though the above assertions are correct, the testcase fails due to the wrong output. This testcase failure log is shown in Fig. 4.
Figure 4: Failure log of full-adder
After correcting the implementation to xor from or, the SVA could pass all the tests, as shown in Fig. 5. This shows that our automated framework can also detect bugs in design implementation.
Figure 5: Success log of full-adder
This shows the usefulness of ChIRAAG in generating SVA from natural language specifications. The basic details of the formatted specification, along with the codebase, simulation files, logs, and prompts, are accessible on our GitHub repository  [^1] Github link will be available upon acceptance.
VI Conclusion
In this work, we have developed ChIRAAG that can generate SV assertion from natural language specification using an LLM as an assertion generator. Primarily, unstructured design specifications are transformed into a structured manner and given to the LLM as an input prompt to generate an SVA suite. The prompting technique is crucial in generating correct assertions. The performance of ChIRAAG on OpenTitan design shows that an LLM is a good starting point to assist an engineer in assertion generation. We noticed only 33% of the generated SVA requires refinement upon iterative prompting. From our observations, it is clear that LLM can be used to produce SV assertion, provided that the LLM will be fed with structured, unambiguous information. We performed our experiment with a gd-LLM; however, we believe a domain-specific LLM for FPV should provide even more accurate assertions in fewer iterations. Although the assertions are correct in satisfying the intended behavior captured by the testbench, it is not possible to confirm 100% functional coverage of design intent. Our future goal will be to check the capability of LLM to address the “consistency” and “completeness” issues in ABV.